OpenStudy (anonymous):

If 3 sin t − cos t = 0, find all possible values of sin t and cos t Holy poo i looked this up but still couldnt find anything that made sense to me

OpenStudy (anonymous):

um, 0?

OpenStudy (anonymous):

Sorry that it isnt it its going to be some sort of fraction and other ugly stuff like square roots

OpenStudy (anonymous):

T.T no one?

OpenStudy (anonymous):

I think it means to solve this equation? 3sint - cost = 0 3sint = cost 3 = cost/sint 1/3 = sint/cost 1/3 = tant arctan(1/3) = t

OpenStudy (anonymous):

well 3 sin - cos = 0 means that 3 sin = cos so that means (3sin)^2 + (sin)^2 = 1 and then you solve theres 4 answers overall